40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 5 21:42:28 2009, Yoichi, Update, PSL, My thoughts on ISS INMONPD_Spectrum_1-10Hz.pdf
    Reply  Fri Feb 6 09:52:35 2009, Kakeru, Update, PSL, Current shunt transfar function TF_CS_gain.pngTF_CS_phase.png
       Reply  Fri Feb 6 23:23:48 2009, Kakeru, Yoichi, Update, PSL, ISS is fixed Current_ShuntTF_gain.pngCurrent_ShuntTF_phase.png
          Reply  Mon Feb 9 16:02:42 2009, Yoichi, Update, PSL, PSL relative intensity noise RIN-13dB.pngRIN-21dB.png
             Reply  Tue Feb 10 23:36:25 2009, Kakeru, Update, PSL, PA current and laser output PA_current_output.png
                Reply  Wed Feb 11 07:28:25 2009, Yoichi, Update, PSL, PA current and laser output 
                   Reply  Wed Feb 11 23:51:53 2009, Kakeru, Update, PSL, PA current and laser output PA_current_output.png
Message ID: 1283     Entry time: Fri Feb 6 23:23:48 2009     In reply to: 1277     Reply to this: 1284
Author: Kakeru, Yoichi 
Type: Update 
Category: PSL 
Subject: ISS is fixed 

Yoichi and me found that the transfar function of the current shunt changed with the current of PA.
We changed PA current and fixed the unstability of ISS.
Now, laser power is stabilized finely, with band of about 1 Hz.
Yoich will post the stabilized noise spectrum.

There looks to be some non-linear relation between PA current and  the TF of current shunt.
It had changed from the TF which we measured yesterday, so it might change again.

I try to write scripts to sweep PA current and measure the laser power and its rms automatically.
It will be apply for auto-adjustment of PA current.


Attached files are the transfar function of the current shunt with changing PA.
They have difference in lower frequency.

Attachment 1: Current_ShuntTF_gain.png  9 kB  | Hide | Hide all | Show all
Current_ShuntTF_gain.png
Attachment 2: Current_ShuntTF_phase.png  8 kB  | Show | Hide all | Show all
ELOG V3.1.3-